An inflated, spherical, beach ball with a radius of 0.3573[m] and average density of 10.65[kg/m3] is being held underwater in a pool by Janelle. The density of the water in the pool is 1000.0[kg/m3]. What is the magnitude of the gravitational force on the ball? Express your answer in N.

I got the answer 1854.72 N but I got it incorrect. Just trying to understand why.
Answer #1

here,

radius of the ball , r = 0.3573 m

the average density of ball, p = 10.65 kg/m^3

the magnitude of the gravitational force on the ball , F = m * g

F = (V * p) * g

F = (4/3 * pi * r^3) * p * g

F = (4/3 * pi * 0.3573^3) * 10.65 * 9.81 N

F = 19.9 N

the magnitude of the gravitational force on the ball is 19.9 N